# Closed Loop Interactive Embodied Reasoning for Robot Manipulation

### v1

@@ C A R Simulation Real
VI Discussion and Conclusions
In this paper, we presented our closed-loop approach for visual and interactive reasoning for robotic manipulation tasks: closed loop interactive embodied reasoning (CLIER). This approach is using our novel MuBlE environment that incorporates MuJoCo physics simulation with high-quality renderer and enables generation of multi-modal demonstration data for robotic manipulation tasks. This fully modular environment enables both data generation and benchmarking of simultaneous reasoning in visual and physical space. CLIER is able to incorporate observations of both visual and physical attributes of the manipulated objects into long-term reasoning. Capturing data from visual and physical measurements in the shared scene graph enables the symbolic reasoning approach and simplifies the implementation of the closed loop approach via keyframe based reasoning. The results from simulated and real-world experiments showed its ability to successfully transfer between the simulated and real environment and to recover from various errors or changes in the scene.
We believe this work provides a new challenge and a benchmark for future systems that aim at bridging the gap between physics simulation for robotic manipulators, realistic visual simulations and execution in real environments.

### v2

@@ C A R Simulation Real
VI Conclusions
In this paper, we presented our Closed-Loop Interactive Embodied Reasoning (CLIER) for robotic manipulation tasks. CLIER is able to incorporate observations of both visual and physical attributes of the manipulated objects into long-term reasoning. Capturing data from visual and physical measurements in the shared scene graph enables the symbolic reasoning approach and simplifies the implementation of the closed loop approach via keyframe-based reasoning. The CLIER utilises MuBlE environment [14] that incorporates MuJoCo physics simulation with high-quality renderer and enables generation of multi-modal demonstration data for robotic manipulation tasks. The results from simulated and real-world experiments showed its ability to successfully transfer between the simulated and real environment and to recover from various errors or changes in the scene.